Find an equation of the line. Write the equation in slope-intercept form.
Through (2,-7); perpendicular to 5x + 2y =18

Answer by algebrapro18(249)

We know that the line we are finding the equation of has a slope that's opposite and reciprical to the slope of the line given because the lines are perpendicular.
So lets find the slope of the line given. To do that we put the equation in slope-intercept form.
5x+2y=18
2y = 18-5x
y = 9 - 5/2x
so we know the slope of the given line is -5/2. So we know the slope of our line is 2/5.
We can now use a formula to find the equation of the line in slope intercept form. We have the point (2,-7) and the slope of the line m = 2/5. So we can use the point-slope formula:
y-y1=m(x-x1)
y-(-7)=2/5(x-2)
y+7=2/5(x-2)
y = 2/5x-4/5-7
y = 2/5x-39/5